A Particle is moving along the circle x2+y2=a2 in anticlockwise direction. The x-y plane is a rough horizontal stationary surface. At the point (acos(th), asin(th)), The unit vector in the direction of friction on the particle is:
  1. coso i^ + sino j^
  2. - [ coso i^ + sino j^]
  3. sino i^ – coso j^
  4. coso i^ – sino j^
Along with proper explaination. 
thanks

Grade:11

1 Answers

Vikas TU
7 years ago
Put (acos(th), asin(th) in the circle eqn. and solve for thetha in differnetial of x2+y2=a
Thus given, x2+y2=a2
2x + 2ydy/dx = 0
at  (acos(th), asin(th)),
Get dy/dx .i.e v and then write a = v^2/a
f = uma
=>